# Reconstruction of Quantum Fields

### v2

@@ I Introduction
which describes distinguishable particles whose single–particle spaces are isomorphic. Indistinguishability is then imposed by projecting onto the symmetric or antisymmetric subspaces, yielding the usual occupation–number description in which only the number of particles per accessible mode is relevant. In this traditional formalism, the exclusivity of Bose and Fermi statistics is assumed rather than derived, and any alternative statistical behaviour is excluded a priori. Nonetheless, a variety of classical and recent works have exhibited consistent generalisations of statistics, including parastatistics Green53; Greenberg1964; HARTLE69; ryan1963; stolt1970; stoilova; WangHazzardProvided, quons Greenberg1991, Gentile statistics Gentile40, and several topological or combinatorial extensions dewitt; myrheim; Wilczek82; Read_2003; fredenhagen89; Surya_2004; Balachandran_2001; Baez06; vanenk2019; ZhangHanssonKivelson1989; Witten1989; Haldane08; Laughlin1983. These are typically introduced by ad hoc modifications of commutation relations, rather than through a structural analysis of indistinguishability.
In this work we develop a systematic treatment of the passage from distinguishable to indistinguishable particles based on operational constraints. We again take as starting point the algebra $T(H)$ whose elements represent labelled multi–particle states, following Segal’s tensor-algebra formulation of second quantization Segal1956a; Segal1956b; Segal1958. Indistinguishability is implemented by an equivalence relation that removes precisely the information operationally inaccessible for distinguishing one constituent from another. The quotient
@@
where $E$ comprises the degrees of freedom accessible to the experimenter (for example the external modes), while $K$ contains internal or hidden degrees of freedom as in treatments of indistinguishable particles with internal structure Benatti2020; Morris2020; NeoriGoyal2012; Neori16; MedinaSanchezDakic2024. On this decomposition we impose three requirements:
@@
Condition 2 has a strong algebraic consequence: it forces $I$ to be quadratically generated. Thus all relations needed to describe indistinguishability arise already at the level of two–particle states, placing our construction within the general theory of quadratic and Koszul algebras Priddy1970; PolishchukPositselski2005; BeilinsonGinzburgSoergel1996; LodayVallette2012; Davydov. This mirrors the bosonic and fermionic cases and contrasts with other generalised statistics whose defining relations may be cubic or of higher order Green53; Greenberg1964; Greenberg1991; Greenberg90; Fivel90; CHEN_1996.
@@
for subspaces $W_{\mathrm{sym}},W_{\mathrm{ext}}\subset K\otimes K$. These internal subspaces are realised as images of projectors on $K\otimes K$, and the existence of an ordered basis (what we called the Poincaré–Birkhoff–Witt property) imposes Yang–Baxter type identities on those projectors, connecting our construction to $R$-matrix and reflection-equation algebras Yang1967; Drinfeld1988; Jimbo1985; Baxter1972; FaddeevReshetikhinTakhtajan1990; Gurevich1991; GurevichPyatovSaponov1996; Majid1993; Majid2000; BM2000; Davydov; BozejkoKummererSpeicher1997
Within this class of quadratic algebras, the single-mode partition functions are exactly those allowed by the operational partition theorem developed by the authors MedinaSanchezDakic2024.
@@
The paper is organised as follows. In Section II we formulate the transition from first to second quantization as a quotient of the tensor algebra and recall Segal’s theorem in this language. Section III introduces the quadratic quotients associated with generalised statistics, proves that order implies quadratic generation, and classifies $\mathrm{U}(d)$–invariant quadratic subspaces. We also derive Yang–Baxter constraints and describe the computation of the single–mode partition function. Section IV develops the corresponding creation and annihilation algebras and their mixed sector, and constructs the transfield representation of $\mathfrak{gl}(d)$. In Section V we work out a concrete finite–order example. Section VI concludes with some remarks and open problems. Technical proofs are collected in an appendix.
@@ II.3 Segal’s theorem
Segal’s theorem shows that, under these invariance and maximality assumptions, the only ideals are the bosonic and fermionic ones.
@@ III Quadratic spaces and generalised statistics
We now introduce the algebraic framework that realises certain transtatistics MedinaSanchezDakic2024 as quadratic quotients of the tensor algebra.
@@ III.4 $\mathrm{U}(d)$ invariance
A $\mathrm{U}(d)$–invariant quadratic subspace $R_{\mathrm{gen}}\subset H\otimes H$ must respect this decomposition. By this reasoning we conclude:
@@
Let $H\cong\mathbb{C}^{d}\otimes K$ with the defining action of $\mathrm{U}(d)$ on $\mathbb{C}^{d}$. Then any $\mathrm{U}(d)$–invariant subspace $R_{\mathrm{gen}}\subset H\otimes H$ is of the form
$R_{\mathrm{gen}}\;=\;\mathrm{Sym}^{2}(\mathbb{C}^{d})\otimes W_{\mathrm{sym}}\;\oplus\;\wedge^{2}(\mathbb{C}^{d})\otimes W_{\mathrm{ext}},$
@@ III.5 Quadratic relations and the generating ideal
$P_{\mathrm{gen}}$ $=P^{(d)}_{\mathrm{sym}}\otimes P^{(K)}_{\mathrm{sym}}+P^{(d)}_{\mathrm{ext}}\otimes P^{(K)}_{\mathrm{ext}}$,
$P_{\mathrm{gen}}^{\perp}$ $=P^{(d)}_{\mathrm{sym}}\otimes P^{(K)\perp}_{\mathrm{sym}}+P^{(d)}_{\mathrm{ext}}\otimes P^{(K)\perp}_{\mathrm{ext}}$.
@@
$H\otimes H\;\cong\;R_{\mathrm{gen}}\;\oplus\;R_{\mathrm{gen}}^{\perp},$
where $R_{\mathrm{gen}}=\mathrm{im}\,P_{\mathrm{gen}}$ is the generating quadratic subspace and $R_{\mathrm{gen}}^{\perp}$ is a chosen complement that will be relevant later to the construction of the annihilation operators.
@@
$r^{\alpha\beta}_{ij}:=P_{\mathrm{gen}}\bigl((e_{i}\otimes e_{j})\otimes(k_{\alpha}\otimes k_{\beta})\bigr).$ (1)
@@
$A:=T(H)/\langle R_{\mathrm{gen}}\rangle.$
@@ III.6 Yang–Baxter constraints and PBW property
Let $P_{\mathrm{gen}}\colon H\otimes H\to R_{\mathrm{gen}}$ be the projector onto the generating subspace. Then the following characterisation is standard in the theory of quadratic algebras.
@@
Let $A=T(H)/\langle R_{\mathrm{gen}}\rangle$ be a quadratic algebra defined by a projector $P_{\mathrm{gen}}$ as above. Then $A$ admits a basis of ordered monomials if and only if, for all $u,v,w\in H$,
$(P_{\mathrm{gen}}\otimes\mathrm{id})(\mathrm{id}\otimes P_{\mathrm{gen}})(P_{\mathrm{gen}}\otimes\mathrm{id})=(\mathrm{id}\otimes P_{\mathrm{gen}})(P_{\mathrm{gen}}\otimes\mathrm{id})(\mathrm{id}\otimes P_{\mathrm{gen}})$ (4)
@@
Proof.
The proof requires elements of the theory of Poincaré-Birkhoff-Witt basis. See appendix. ∎
The identity (4) is a quantum Yang–Baxter type equation. When we insert the explicit form of $P_{\mathrm{gen}}$ in terms of symmetric and antisymmetric external projectors and internal projectors $P^{(K)}_{\mathrm{sym}}$ and $P^{(K)}_{\mathrm{ext}}$, the condition decouples into two internal Yang–Baxter identities
@@ III.7 Single–mode partition function
The coefficients $g_{n}$ are determined solely by the quadratic projector $P_{\mathrm{gen}}$ acting on $K\otimes K$. For each $n\geq 0$ set $V_{n}=K^{\otimes n}$ and define operators
$P^{(n)}_{k}:=\mathrm{id}^{\otimes(k-1)}\otimes P_{\mathrm{gen}}\otimes\mathrm{id}^{\otimes(n-k-1)},\qquad 1\leq k\leq n-1.$
@@
In the language of MedinaSanchezDakic2024 we know that the Hilbert-Poincaré series of the algebra as defined here is in closely related to the partition function, once we exchange the variable $t$ with the Boltzmann factor $e^{\beta\,\epsilon}$. We can then finally state the main theorem:
Theorem 2 (Quadratic realisation of transtatistics).
Let $H\cong\mathbb{C}^{d}\otimes K$ be a finite dimensional one–particle Hilbert space, where $\mathbb{C}^{d}$ carries the operational $U(d)$ action on accessible modes and $K$ encodes internal degeneracies. Let $T(H)=\bigoplus_{n\geq 0}H^{\otimes n}$ be the tensor algebra and let $I\subset T(H)$ be a homogeneous ideal such that the Fock space
F:=T(H)/IF\;:=\;T(H)/I
satisfies the following properties:
@@
    PBW property. There exists a total order on a homogeneous basis $\{X_{i\alpha}\}$ of $H$ such that the equivalence classes of ordered monomials in the $X_{i\alpha}$ form a linear basis of $F$.
@@
    $U(d)$ invariance. The accessible unitary group $U(d)$ acts on $H$ via $u\cdot(e_{i}\otimes f_{\alpha})=(ue_{i})\otimes f_{\alpha}$ and this action extends to $T(H)$ leaving $I$ invariant.
Then the following holds. The single–mode generating function $G(t)$ is the character of an elementary transtatistics MedinaSanchezDakic2024. Equivalently, there exist integral polynomials $Q_{\pm}(t)$ with all roots real and of fixed sign,
$Q_{+}(t)=\prod_{i}(1-\alpha_{i}t),\quad Q_{-}(t)=c_{-}\prod_{i}(1+\beta_{i}t),$
for $\alpha_{i},\beta_{i}>0,\,c_{-}\in\mathbb{N}$, such that
$G(t)\;=\;Q_{-}(t)\quad\text{or}\quad G(t)\;=\;\frac{1}{Q_{+}(t)}.$
In this case we say that $F$ realises a transfermionic statistics of type $q_{0},q_{1},\dots]_{-}$ or a transbosonic statistics of type $q_{0},q_{1},\dots]_{+}$ respectively.
Proof.
See appendix. ∎
This shows that we can generate the partition functions presented in MedinaSanchezDakic2024 using quadratically generated algebras and more generally that at least for some of them there exists a first quantization picture that can be lifted to a second quantization one. Interestingly, a classical result due to Priddy shows that algebras admitting a Poincaré–Birkhoff–Witt (PBW) basis are Koszul Priddy1970. This implies the existence of a precise duality between the Hilbert–Poincaré series of the algebra and that of its Koszul dual. Concretely, if

### v3

@@ I Introduction
which describes distinguishable particles whose single–particle spaces are isomorphic. Indistinguishability is then imposed by projecting onto the symmetric or antisymmetric subspaces, yielding the usual occupation–number description in which only the number of particles per accessible mode is relevant. In this traditional formalism, the exclusivity of Bose and Fermi statistics is assumed rather than derived, and any alternative statistical behaviour is excluded a priori. Nonetheless, a variety of classical and recent works have exhibited consistent generalisations of statistics, including parastatistics Green (1953); Greenberg (1964); HARTLE and TAYLOR (1969); Ryan and Sudarshan (1963); Stolt and Taylor (1970); Stoilova and Van der Jeugt (2020); Wang and Hazzard (2025), quons Greenberg (1991), Gentile statistics Gentile j. (1940), and several topological or combinatorial extensions Laidlaw and DeWitt (1971); Leinaas and Myrheim (1977); Wilczek (1982); Read (2003); Fredenhagen et al. (1989); Surya (2004); Balachandran et al. (2001); Baez et al. (2006); van Enk (2019); Zhang et al. (1989); Witten (1989); Bernevig and Haldane (2008); Laughlin (1983). These are typically introduced by ad hoc modifications of commutation relations, rather than through a structural analysis of indistinguishability.
In this work we develop a systematic treatment of the passage from distinguishable to indistinguishable particles based on operational constraints. We again take as starting point the algebra $T(H)$ whose elements represent labelled multi–particle states, following Segal’s tensor-algebra formulation of second quantization Segal (1956a, b, 1958). Indistinguishability is implemented by an equivalence relation that removes precisely the information operationally inaccessible for distinguishing one constituent from another. The quotient
@@
where $E$ comprises the degrees of freedom accessible to the experimenter (for example the external modes), while $K$ contains internal or hidden degrees of freedom as in treatments of indistinguishable particles with internal structure Benatti et al. (2020); Morris et al. (2020); Neori and Goyal (2012); Neori (2016); Sánchez and Dakić (2024). On this decomposition we impose three requirements:
@@
Condition 2 has a strong algebraic consequence: it forces $I$ to be quadratically generated. Thus all relations needed to describe indistinguishability arise already at the level of two–particle states, placing our construction within the general theory of quadratic and Koszul algebras Priddy (1970); Polishchuk and Positselski (2005); Beilinson et al. (1996); Loday and Vallette (2012); Davydov (2000). This mirrors the bosonic and fermionic cases and contrasts with other generalised statistics whose defining relations may be cubic or of higher order Green (1953); Greenberg (1964, 1991, 1990); Fivel (1990); CHEN et al. (1996).
@@
for subspaces $W_{\mathrm{sym}},W_{\mathrm{ext}}\subset K\otimes K$. These internal subspaces are realised as images of projectors on $K\otimes K$, and the existence of an ordered basis (what we called the Poincaré–Birkhoff–Witt property) imposes Yang–Baxter type identities on those projectors, connecting our construction to $R$-matrix and reflection-equation algebras Yang (1967); Drinfel’d (1988); Jimbo (1985); Baxter (1972); Faddeev et al. (1990); Gurevich (1991); Gurevich et al. (1996); Majid (1993, 2000); Borowiec and Marcinek (2000); Davydov (2000); Bożejko et al. (1997)
Within this class of quadratic algebras, which we call PBW $U(d)$-equivariant algebras, we obtain a complete characterisation of the admissible single-mode partition functions. Our main result (Theorem 2 or Field Theorem) states that a formal power series $G(t)\in\mathbb{Z}_{\geq 0}[[t]]$ arises as the single-mode Hilbert–Poincaré series of such an algebra if and only if it is a rational function of the form
$G(t)=\frac{Q_{-}(t)}{Q_{+}(t)},$
where $Q_{\pm}(t)\in\mathbb{Z}[t]$, $Q_{+}(0)=1$, and all roots of $Q_{+}$ (resp. $Q_{-}$) are real and strictly positive (resp. strictly negative). This coincides precisely with the class of partition functions identified by the operational Partition Theorem developed in Sánchez and Dakić (2024).
@@
The paper is organised as follows. In Section II we formulate the transition from first to second quantization as a quotient of the tensor algebra and recall Segal’s theorem in this language. Section III introduces the quadratic quotients associated with generalised statistics, proves that order implies quadratic generation, and classifies $\mathrm{U}(d)$–invariant quadratic subspaces. We also derive Yang–Baxter constraints and describe the computation of the single–mode partition function. Section IV develops the corresponding creation and annihilation algebras and their mixed sector, and constructs the transfield representation of $\mathfrak{gl}(d)$. In Section V we work out a concrete finite–order example. Section VI concludes with some remarks and open problems. All proofs are collected in the appendix.
@@ II.3 Segal’s theorem
Segal’s theorem draws a sharp boundary: if one insists on maximality and full invariance of the ideal under the creation operators themselves, the only possible statistics are Bose and Fermi. This is a powerful rigidity result, but its strength comes precisely from the stringency of its hypotheses. In particular, the maximality condition privileges the largest possible identification of states, collapsing the ideal to one of two extremes. If instead one is willing to consider ideals that are invariant under mode transformations but not necessarily maximal, a much richer landscape of consistent statistics opens up.
The question then becomes structural: among all homogeneous ideals compatible with the operational requirements laid out in the introduction, which ones lead to well-defined many-particle spaces with a sensible physical interpretation? The answer, as we now show, is that the three conditions imposed on the Fock space (a graded particle-number decomposition, an ordered monomial basis, and invariance under unitary mode transformations) are not only physically natural but also algebraically decisive. Together they force the ideal to be quadratically generated, placing the entire construction within the well-developed framework of quadratic and Koszul algebras
@@ III Quadratic spaces and generalised statistics
We now introduce the algebraic framework that realises certain transtatistics Sánchez and Dakić (2024) as quadratic quotients of the tensor algebra.
@@ III.1 Internal degeneracies and quantum grammar
The restriction to quadratic relations admits a direct conceptual justification in terms of indistinguishability. In the standard formulation, indistinguishability is implemented by requiring invariance of the state under the action of the symmetric group, whose elements are generated by pairwise transpositions. Thus, the indistinguishability of an arbitrary number of particles is completely determined by the consistency of exchanges at the level of particle pairs. In the present framework, this principle is reflected algebraically by requiring that all relations encoding indistinguishability arise from constraints in the two-particle sector. Higher-degree relations would introduce additional structure not dictated by pairwise exchange, and would therefore go beyond the minimal requirement of indistinguishability. In this sense, quadratic generation captures precisely the idea that indistinguishability is a pairwise notion, and ensures that the extension to many-particle states is governed entirely by consistent two-body exchanges.
@@ III.4 $\mathrm{U}(d)$ invariance
A $\mathrm{U}(d)$–invariant quadratic subspace $R_{\mathrm{g}}\subset H\otimes H$ must respect this decomposition. By this reasoning we conclude:
@@
Let $H\cong\mathbb{C}^{d}\otimes K$ with the defining action of $\mathrm{U}(d)$ on $\mathbb{C}^{d}$. Then any $\mathrm{U}(d)$–invariant subspace $R_{\mathrm{g}}\subset H\otimes H$ is of the form
$R_{\mathrm{g}}\;=\;\mathrm{Sym}^{2}(\mathbb{C}^{d})\otimes W_{\mathrm{sym}}\;\oplus\;\wedge^{2}(\mathbb{C}^{d})\otimes W_{\mathrm{ext}},$
@@ III.5 Quadratic relations and the generating ideal
$P_{\mathrm{g}}$ $=P^{(d)}_{\mathrm{sym}}\otimes P^{(K)}_{\mathrm{sym}}+P^{(d)}_{\mathrm{ext}}\otimes P^{(K)}_{\mathrm{ext}}$,
$P_{\mathrm{g}}^{\perp}$ $=P^{(d)}_{\mathrm{sym}}\otimes P^{(K)\perp}_{\mathrm{sym}}+P^{(d)}_{\mathrm{ext}}\otimes P^{(K)\perp}_{\mathrm{ext}}$.
@@
$H\otimes H\;\cong\;R_{\mathrm{g}}\;\oplus\;R_{\mathrm{g}}^{\perp},$
where $R_{\mathrm{g}}=\mathrm{im}\,P_{\mathrm{g}}$ is the generating quadratic subspace and $R_{\mathrm{g}}^{\perp}$ is a chosen complement that will be relevant later to the construction of the annihilation operators.
@@
$r^{\alpha\beta}_{ij}:=P_{\mathrm{g}}\bigl((e_{i}\otimes e_{j})\otimes(k_{\alpha}\otimes k_{\beta})\bigr).$ (1)
@@
$A:=T(H)/\langle R_{\mathrm{g}}\rangle.$
@@ III.6 Yang–Baxter constraints and PBW property
Let $P_{\mathrm{g}}\colon H\otimes H\to R_{\mathrm{g}}$ be the projector onto the generating subspace. Then the following characterisation is standard in the theory of quadratic algebras.
@@
Let $A=T(H)/\langle R_{\mathrm{g}}\rangle$ be a quadratic algebra defined by a projector $P_{\mathrm{g}}$ as above. Then $A$ admits a basis of ordered monomials if and only if, for all $u,v,w\in H$,
$(P_{\mathrm{g}}\otimes\mathrm{id})(\mathrm{id}\otimes P_{\mathrm{g}})(P_{\mathrm{g}}\otimes\mathrm{id})=(\mathrm{id}\otimes P_{\mathrm{g}})(P_{\mathrm{g}}\otimes\mathrm{id})(\mathrm{id}\otimes P_{\mathrm{g}})$ (4)
@@
The identity (4) is a quantum Yang–Baxter type equation. When we insert the explicit form of $P_{\mathrm{g}}$ in terms of symmetric and antisymmetric external projectors and internal projectors $P^{(K)}_{\mathrm{sym}}$ and $P^{(K)}_{\mathrm{ext}}$, the condition decouples into two internal Yang–Baxter identities
@@ III.7 Single–mode partition function
The coefficients $g_{n}$ are determined solely by the quadratic projector $P_{\mathrm{g}}$ acting on $K\otimes K$. For each $n\geq 0$ set $V_{n}=K^{\otimes n}$ and define operators
$P^{(n)}_{k}:=\mathrm{id}^{\otimes(k-1)}\otimes P_{\mathrm{g}}\otimes\mathrm{id}^{\otimes(n-k-1)},\qquad 1\leq k\leq n-1.$
@@
In the language of Sánchez and Dakić (2024) we know that the Hilbert-Poincaré series of the algebra as defined here is in closely related to the partition function, once we exchange the variable $t$ with the Boltzmann factor $e^{\beta\,\epsilon}$. We can then finally state the main theorem:
Theorem 2 (Field).
Let $G(t)\in\mathbb{Z}_{\geq 0}[[t]]$. The following are equivalent:
@@
    There exists a homogeneous quadratic quotient
    $A=T(H)/\langle R\rangle,\qquad H\simeq\mathbb{C}^{d}\otimes K,$
    which is $U(d)$-invariant and admits a PBW basis, such that the associated single-mode Hilbert–Poincaré series is $G(t)$.
@@
    $G(t)$ is a rational function of the form
    $G(t)=\frac{Q_{-}(t)}{Q_{+}(t)},$
    where $Q_{\pm}(t)\in\mathbb{Z}[t]$, $Q_{+}(0)=1$, and all roots of $Q_{+}$ (resp. $Q_{-}$) are real and strictly positive (resp. strictly negative).
This shows that we can generate the partition functions presented in Sánchez and Dakić (2024) using quadratically generated algebras and more generally that there exists a first quantization picture that can be lifted to a second quantization one. Interestingly, a classical result due to Priddy shows that algebras admitting a Poincaré–Birkhoff–Witt (PBW) basis are Koszul Priddy (1970). This implies the existence of a precise duality between the Hilbert–Poincaré series of the algebra and that of its Koszul dual. Concretely, if
@@
In a complementary direction, one has from the canonical theory of PBW algebras that the Hilbert–Poincaré series is necessarily rational. More precisely, by Corollary 6.2 Polishchuk and Positselski (2005), for any subset $S\subseteq[1,m]^{2}$, let $M_{S}$ denote the $m\times m$ matrix with entries
$(M_{S})_{ij}=\begin{cases}1&\text{if }(i,j)\in S,\\ 0&\text{otherwise},\end{cases}$
and let $\bar{S}=[1,m]^{2}\setminus S$ be its complement, with associated matrix $M_{\bar{S}}$. Then the Hilbert series of a PBW algebra $A$ takes the form
$h_{A}(z)=\frac{\det(1+zM_{\bar{S}})}{\det(1-zM_{S})}.$
We conclude that for every generating function $G(t)$ arising in Theorem 2, there exists an integer $m$ and a subset $S\subseteq[1,m]^{2}$ such that
$G(t)=\frac{\det(1+tM_{\bar{S}})}{\det(1-tM_{S})}.$
Thus, the admissible class of Hilbert–Poincaré series is exactly realised by ratios of characteristic determinants of 011 matrices associated with complementary subsets of $1,m]^{2}$. In particular, the analytic structure of $G(t)$ is encoded by the spectral data of the pair $(M_{S},M_{\bar{S}})$, providing a concrete combinatorial realisation of the abstract classification given by Theorem 2.
It is worth placing this result as well in the context of classical results on total positivity. A sequence $(a_{n})_{n\geq 0}$ is a Pólya frequency (PF) sequence if and only if its associated Toeplitz matrix is totally nonnegative, that is, if and only if it is totally positive. In Reiner and Welker (2005), the authors ask for which Koszul algebras the Hilbert function is a PF sequence. Within our framework, Theorem 2 provides an explicit construction of a large class of such sequences: the single-mode Hilbert–Poincaré series arising from PBW $U(d)$-equivariant algebras are precisely of the form $G(t)=Q_{-}(t)/Q_{+}(t)$ with $Q_{\pm}$ having real roots of fixed sign, and therefore generate totally positive (PF) coefficient sequences of the rational type. In this sense, our result realises a distinguished subclass of PF Hilbert functions through a concrete algebraic and physically meaningful construction keeping in mind that these series by Sánchez and Dakić (2024) are in one-to-one correspondence with partition functions. In that sense we have stablished a complete formal classification of the link between partition functions, state spaces and as we will see in the next section, creation-annihilation algebras.